OpenStudy (asad786):

WILL MEDAL Which expression is equivalent to this one? 9(x – 23) A. 9x – 23 B. 9x – 9 · 23 C. 9x · (–23) D. 9 + x – 23

OpenStudy (swamijatin2):

no you understand by that example .. now you are able to do

OpenStudy (anonymous):

Not A, Not D. What do you think?

OpenStudy (swamijatin2):

1st u tell your anwer

OpenStudy (asad786):

b

OpenStudy (welshfella):

a(b - c) = ab - ac

OpenStudy (asad786):

what

OpenStudy (igreen):

B is correct..

OpenStudy (welshfella):

this is called the Distributive Law

OpenStudy (asad786):

oh nvm its c

OpenStudy (swamijatin2):

yes u r right

OpenStudy (asad786):

wait so i was right about b

OpenStudy (igreen):

No..you were right in the beginning.

OpenStudy (asad786):

oh thanks

OpenStudy (welshfella):

not c B is correct

OpenStudy (igreen):

\(\sf 9(x - 23) \rightarrow (9 \cdot x) + (9 \cdot -23) \rightarrow 9x + 9 \cdot -23\)

OpenStudy (anonymous):

B is correct.

OpenStudy (igreen):

\(\sf 9 \cdot -23\) is the same as \(\sf -9 \cdot 23\)